Didn't really understand the problem, its says:
Which of the following gives the fractions 2/3, 5/8, and 7/11 arranged in order from least to greatest?

Found 2 solutions by drk, jim_thompson5910:
Answer by drk(1908)

To arrange fractions from least to greatest you have two options:(1) compare them two at a time or (2) find a common denominator. Our fractions were:
2/3, 5/8, and 7/11.
common denominator is 3*8*11.
Once you have common denominator, change the numerators. We get: the fractions become 2(88)/3*8*11 , 5(33)/3*8*11, 7(21)/3*8*11.
Now 2*88 = 176, 5*33 = 165 and 7*21 = 147.
In order from least to greatest, we then get:
7/11 , 5/8, 2/3.

Answer by jim_thompson5910(35256)

The order should be 5/8, 7/11, 2/3 since 5/8 is smaller than 7/11